OpenStudy (anonymous):

Let y=(6 x+1/x)^(-2). Then (dy)/(dx) = A(6 x+1/x)^B (6-1/(x^C))^D, where

OpenStudy (anonymous):

how can I find A,B,C, & D

OpenStudy (anonymous):

chain rule in action

OpenStudy (anonymous):

\[y=(6x+\frac{1}{x})^{-2}\] \[y'=-2\times (6x+\frac{1}{x})^{-3}\times \frac{d}{dx}[(6x+\frac{1}{x})\] \[y'=-2(x6+\frac{1}{x})^{-3}(6-\frac{1}{x^2})\]